Airline fare for a flight from elmwood to Palmdale are $85 for first class and $60 for coach class.on Friday,79 passengers paid a total of 5290 to fly from elmwood to Palmdale .How many of each type of ticket were sold?

Answer by mananth(16946)

x First Class
y Coach class
Total passengers
1 x + 1 y = 79 .............1
Total fare for all passengers
85 x + 60 y = 5290.00 .............2
Eliminate y
multiply (1)by -60
Multiply (2) by 1
-60 x -60 y = -4740
85 x + 60 y = 5290
Add the two equations
25 x = 550
/ 25
x = 22
plug value of x in (1)
1 x + 1 y = 79
22 + y = 79
y = 79 -22
y = 57

22 First Class passengers
57 Coach class passengers
m.ananth@hotmail.ca